# FinAgentBench: A Benchmark Dataset for Agentic Retrieval in Financial Question Answering

### v2

@@ Abstract.
Accurate information retrieval (IR) is critical in the financial domain, where investors must identify relevant information from large collections of documents. Traditional IR methods—whether sparse or dense—often fall short in retrieval accuracy, as it requires not only capturing semantic similarity but also performing fine-grained reasoning over document structure and domain-specific knowledge. Recent advances in large language models (LLMs) have opened up new opportunities for retrieval with multi-step reasoning, where the model ranks passages through iterative reasoning about which information is most relevant to a given query. However, there exists no benchmark to evaluate such capabilities in the financial domain. To address this gap, we introduce FinAgentBench, the first large-scale benchmark for evaluating retrieval with multi-step reasoning in finance – a setting we term agentic retrieval. The benchmark consists of 3,429 expert-annotated examples on S&P-100 listed firms and assesses whether LLM agents can (1) identify the most relevant document type among candidates, and (2) pinpoint the key passage within the selected document. Our evaluation framework explicitly separates these two reasoning steps to address context limitations. This design enables to provide a quantitative basis for understanding retrieval-centric LLM behavior in finance. We evaluate a suite of state-of-the-art models and further demonstrated how targeted fine-tuning can significantly improve agentic retrieval performance. Our benchmark provides a foundation for studying retrieval-centric LLM behavior in complex, domain-specific tasks for finance. We will release the dataset publicly upon acceptance of the paper and plan to expand and share dataset for the full S&P 500 and beyond.
@@ 1. Introduction
To address this gap, we introduce the first large-scale benchmark, FinAgentBench, for evaluating generative retrieval systems in the financial domain, focusing on a setting we term agentic retrieval. Unlike prior benchmarks that assess retrieval in a single stage, our proposed benchmark evaluates the ability of LLM agents to reason and retrieve the relevant information through a two-stage pipeline: (1) identifying which document to retrieve, evaluated via document-level ranking, and (2) determining which part of the document to focus on, evaluated via chunk-level ranking. The benchmark contains 3,429 samples and is constructed from real-world financial documents paired with expert-written queries and annotations, reflecting authentic use cases faced by professional investors. By capturing both retrieval accuracy and reasoning depth, it provides a foundation for systematically analyzing the strengths and limitations of LLM-based generative retrieval in high-stakes domains like finance.
@@
  • We propose FinAgentBench, to the best of our knowledge the first large-scale benchmark for evaluating agentic retrieval in finance, featuring 3,429 expert-annotated samples across document- and chunk-level ranking tasks.
@@ 3.2. Data Collection
In this paper, we curate a subset of the corpus by focusing exclusively on companies listed in the S&P-100 (OEX) index. This filtering yields a more targeted dataset comprising approximately 500 documents, which serves as the primary scope of analysis throughout the study. We initially start with the S&P-100 as a first set, with plans to expand and share results for the full S&P 500 and beyond.
@@ 3.3. Data Annotation
All annotations were conducted by finance professionals with relevant industry experience. Each query was independently labeled by two annotators, and disagreements were resolved through a brief adjudication process. This dual-annotator setup with adjudication effectively functions as a cross-validation mechanism, enhancing the reliability and consistency of the labels. The final dataset comprises 5,00 annotated examples for document-type ranking and 2,929 fully annotated queries for chunk-level relevance, covering 10 U.S. public firms listed in the S&P-100 (OEX) index. On average, each query is associated with 2.92 relevant document types, 6.41 directly relevant chunks, and 9.24 partially relevant chunks.
@@ 4.1. Experimental Setup
Performance is measured using standard ranking metrics measure by top-$5$ results: normalized Discounted Cumulative Gain (nDCG), Mean Average Precision (MAP), and Mean Reciprocal Rank (MRR). For chunk-level evaluation, we report metrics over the top ranked passages, evaluated against expert-annotated graded relevance scores. All models operate in a retrieval-only setting without access to external tools or retrieval augmentation.

### v3

@@ Abstract.
Accurate information retrieval (IR) is critical in the financial domain, where investors must identify relevant information from large collections of documents. Traditional IR methods—whether sparse or dense—often fall short in retrieval accuracy, as it requires not only capturing semantic similarity but also performing fine-grained reasoning over document structure and domain-specific knowledge. Recent advances in large language models (LLMs) have opened up new opportunities for retrieval with multi-step reasoning, where the model ranks passages through iterative reasoning about which information is most relevant to a given query. However, there exists no benchmark to evaluate such capabilities in the financial domain. To address this gap, we introduce FinAgentBench, the first large-scale benchmark for evaluating retrieval with multi-step reasoning in finance – a setting we term agentic retrieval. The benchmark consists of 26K expert-annotated examples on S&P-500 listed firms and assesses whether LLM agents can (1) identify the most relevant document type among candidates, and (2) pinpoint the key passage within the selected document. Our evaluation framework explicitly separates these two reasoning steps to address context limitations. This design enables to provide a quantitative basis for understanding retrieval-centric LLM behavior in finance. We evaluate a suite of state-of-the-art models and further demonstrated how targeted fine-tuning can significantly improve agentic retrieval performance. Our benchmark provides a foundation for studying retrieval-centric LLM behavior in complex, domain-specific tasks for finance.
@@ 1. Introduction
To address this gap, we introduce the first large-scale benchmark, FinAgentBench, for evaluating generative retrieval systems in the financial domain, focusing on a setting we term agentic retrieval. Unlike prior benchmarks that assess retrieval in a single stage, our proposed benchmark evaluates the ability of LLM agents to reason and retrieve the relevant information through a two-stage pipeline: (1) identifying which document to retrieve, evaluated via document-level ranking, and (2) determining which part of the document to focus on, evaluated via chunk-level ranking. The benchmark contains about 26K samples and is constructed from real-world financial documents paired with expert-written queries and annotations, reflecting authentic use cases faced by professional investors. By capturing both retrieval accuracy and reasoning depth, it provides a foundation for systematically analyzing the strengths and limitations of LLM-based generative retrieval in high-stakes domains like finance.
@@
  • We propose FinAgentBench, to the best of our knowledge the first large-scale benchmark for evaluating agentic retrieval in finance, featuring about 26K expert-annotated samples across document- and chunk-level ranking tasks.
@@ 3.2. Data Collection
In this paper, we curate a subset of the corpus by focusing exclusively on companies listed in the S&P-500 index. This filtering yields a more targeted dataset comprising approximately 500 documents, which serves as the primary scope of analysis throughout the study.
@@
All annotations were conducted by finance professionals with relevant industry experience. Each query was independently labeled by two annotators, and disagreements were resolved through a brief adjudication process. This dual-annotator setup with adjudication effectively functions as a cross-validation mechanism, enhancing the reliability and consistency of the labels.
@@ 4.1. Experimental Setup
Performance is measured using standard ranking metrics measure by top-55 results: normalized Discounted Cumulative Gain (nDCG), Mean Average Precision (MAP), and Mean Reciprocal Rank (MRR). For chunk-level evaluation, we report metrics over the top ranked passages, evaluated against expert-annotated graded relevance scores. All models operate in a retrieval-only setting without access to external tools or retrieval augmentation.